These were accepted.
Stephen will continue to send out the report with the names of the machines requiring rebooting. Ongoing - but urgent for machines with firewall holes.
Labels have been added to the spare machines in the Forum server room. George reminded us to update the appropriate fibre wiki page when moving kit.
An account has now been created and tested. Iain has sent out details to students with the caveat that IS will not provide support for IMAP.
Blog articles suggested:
Report from Computing Executive Group
Reports from units
Please update the wiki page to include any services that you think might be affected.
This was discussed at length. In addition to the sanity checks that George has already added, it was suggested that regular checks should be made to ensure that the firewall is in place e.g. by checking ports that we know should never be open. Alastair also commented that we should try to limit what we do just before extended breaks although this was a genuine seemingly innocuous typo.
We clarified that should a VM need to continue running while the KVM server is undergoing maintenance, the criticality should be set to HIGH. MPU will then automatically migrate the service.
IMAP essentially works but will not be supported by IS. Iain has emailed students with details. It would appear that some students never received notification about the move to Office365. The recommendation is that we should ask a number of our students when they return for semester 2. We can then decide on any appropriate action.
See also Infrastructure report. It was suggested that an email to sys-announce explaining what happened would be appropriate. Neil sent this shortly after the meeting finished. It points to a blog article for further details.
There were no reports but it was agreed that groups should circulate their choices and decide which they are planning to take further to avoid any duplication.
Topics for discussion
It was agreed that this would only be useful on machines with multiple user logins (e.g. ssh, login and compute servers) and machines running a web service. It can otherwise be problematic e.g. with AFS servers.
The next meeting will be on Wednesday 23rd January at 10am in IF-4.31.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh